# VisionUnite: A Vision-Language Foundation Model for Ophthalmology Enhanced with Clinical Knowledge

### v1

@@ Methods
The architecture of VisionUnite To enable the model to achieve open-ended prediction and have user interaction and clinical interpretation capabilities, we propose a large vision-language foundation model with clinical knowledge. Illustrated in 1 (c), our proposed VisionUnite comprises a Transformer-based vision encoder, a vision adapter tailored for signs-level classification, a vision projector for aligning visual embeddings and text tokens, and a substantial language model fine-tuned on llama-7B. The large language model is different from the vision model in that it can achieve open-ended prediction. By leveraging the combination of image-text pairs and extensive dialogue information, VisionUnite excels in responding to user questions based on visual data. We introduce visual features with sign-level features into the model, which improve its ability to understand images and further enhance its clinical interpretation ability. This design can extend its ability to analyze fundus images for medical diagnosis. Our model is also designed for multi-round dialogues, which helps the model follow user instructions and achieve better problem-understanding ability. Harnessing the synergy of image-text pairs and multi-round dialogue information, VisionUnite demonstrates proficiency in addressing user queries grounded in visual data, extending to the capability of diagnosing fundus images. The holistic architecture of VisionUnite enables nuanced and contextually informed responses, ensuring a comprehensive understanding of the inquiries based on visual elements. We leverage the EVA02 model [25] with the CLIP Vision Encoder to collectively serve as the Transformer-Based Vision Encoder. Our configuration incorporates 12 layers of EvaBlock [25] within the original vision Transformer Block [26]. Notably, in our design, the GELU activation [27] in the original vision Transformer Block is replaced with the more advanced SwiGLU [28]. It is also widely used in PaLM [29], LLaMA [11], and LLaMA-Adapter [30] to enhance the performance of the FFN (position feedforward network) layer in the Transformer architecture. Specifically, images, text descriptions, and questions are fed into corresponding encoders to form visual embeddings, text embeddings, and dialogue tokens during the training phase. In this phase, we use contrastive learning to supervise the encoding formation of visual embeddings and text embeddings. Afterward, we use the vision adapter to perform secondary encoding and sign classification on visual embeddings, which form six corresponding sign category embeddings. Then we concatenate the visual embeddings and sign category embeddings together as inputs to the vision projector. For the text features, we embed the corresponding sign categories ($CLS$ tokens) and concatenate them to form text features with the dialogue tokens based on the prediction of the signs.
The vision encoder in our VisionUnite model facilitates the extraction of visual embeddings, serving as input for the vision adapter to generate classification predictions at the sign level. The prediction of signs by VisionUnite plays a pivotal role in improving both diagnostic accuracy and efficiency. By accurately identifying and localizing signs such as lesions in the images, the model narrows the diagnostic possibilities and guides the model toward more targeted investigations. It not only optimizes the diagnostic process by reducing the need for too broad differential diagnoses but also enhances accuracy by focusing on the most relevant clinical signs. In addition, the model is designed with the flexibility to accommodate an expanding array of sign categories, enhancing its adaptability for future developments. To further refine VisionUnite, we plan to increase the universality of the model by supporting a wider range of sign-level labels. The expansion is aimed at improving the granularity of diagnostic evidence and enhancing the interpretability of the outputs from the models. VisionUnite can recognize six primary signs, which cover a wide range of ophthalmic diseases. This kind of capability aligns closely with the diagnostic process used by medical professionals, who often start their diagnosis by identifying specific signs before narrowing down to a more precise disease identification. Furthermore, the ability of VisionUnite to leverage these predicted image signs effectively distinguishes it from other ophthalmic deep learning models. By focusing on sign-level features, VisionUnite ensures that the diagnostic suggestions it provides are both precise and clinically relevant, offering a robust tool that supports clinicians in making informed decisions. Following the acquisition of predicted sign-level data, VisionUnite seamlessly employs the vision projector to synchronize visual embeddings with dialogue tokens. It entails concatenating the visual embedding with the class embedding, which is then fed into the vision projector. The alignment process involves matching the resulting feature dimensions with $CLS$ (class) and dialogue tokens via concatenation, with the $CLS$ token encapsulating the predicted sign-level feature. The multimodal features, which combine visual embeddings, $CLS$ tokens, and dialogue tokens, are the inputs for training the final text generation response within the framework of the large language model. Overall, the integration of state-of-the-art components contributes to the overall efficacy and robustness of VisionUnite, positioning it as a cutting-edge vision-language model.
@@
The training objectives of VisionUnite In the architecture of our proposed VisionUnite, we have crafted three distinct training objectives to enhance convergence: image-text contrastive learning, classification supervised learning, and text-generation supervised learning. The utilization of image-text contrastive learning facilitates the refinement of visual encoders, aiding them in more effectively aligning fundus image features. Meanwhile, the application of classification-supervised learning contributes by furnishing accurate feature categories for both visual embeddings and dialogue tokens. We utilize the accuracy to guide the model training process, enhancing its overall performance. Finally, text-generation supervised learning plays a role in guiding the output of the language model, which is pivotal for achieving accurate and open-ended disease diagnoses. Unlike previous vision models, large language models are trained on vast amounts of diverse textual data. The extensive textual data enables them to generate context-aware responses and detailed explanations that are coherent and clinically relevant. We can achieve open-ended disease diagnosis using the generated text from language models. In contrast, the prediction categories of vision models are limited and concentrated on specific diseases. Collectively, these three training objectives synergistically contribute to the robust convergence and performance of VisionUnite, making it a versatile and effective tool in the field of ophthalmic diagnostics.
@@
Image-text contrastive learning: To attain seamless alignment between image and text features, we employ image-text contrastive learning. In particular, we leverage the CLIP loss methodology to quantify the similarity between image embeddings and text embeddings, as illustrated in Figure 1 (d) and below:
@@
Classification supervised learning: In facilitating the acquisition of sign-level features, we employ conventional classification learning to guide the training of VisionUnite. Notably, we embrace multi-label classification while acknowledging that each sample may encompass more than one predicted category. This approach aligns more closely with the intricacies of real-world clinical scenarios, where patients may concurrently experience multiple types of eye diseases. For each category under supervision, we apply cross-entropy loss, aggregating these individual losses to derive the ultimate classification loss, as depicted in Figure 1 (e) and below:
$L_{CLS}=\sum_{k=1}^{M}{L_{CLS,k}}=-\frac{1}{N}\sum_{k=1}^{M}{\sum_{i=1}^{N}[y_{k,i}\cdot log(p_{k,i})+(1-y_{k,i})\cdot log(1-p_{k,i})]}$ (4)
where $M$ and $N$ denote the number of categories and samples respectively. $L_{CLS,k}$ signifies the cross-entropy loss for category $K$. And $y_{k,i}$ denotes whether sample $i$ belongs to class $K$, and $p_{k,i}$ represents the probability of sample $i$ belonging to class $K$.
Text-generation supervised learning: Within the framework of VisionUnite, we employ text-generation supervised learning to guide the text output of LLM, which is an essential aspect given that the generated text must intuitively articulate the diagnostic results and their underlying rationale. The LLM loss aims to train the model to generate text closely resembling patterns in its training data as shown in Figure 1 (f), and its formulation is articulated as follows:
@@
Datasets for pre-training VisionUnite
In the pre-training phase of VisionUnite, we utilize a dataset comprising 1.24 million image-text pairs, encompassing around 616,435 natural image-text pairs sourced from the COCO dataset [12] and around 623,816 biomedical image-text pairs sourced from the PMC-OA dataset [13, 14] (608,027 images) and Retina Image Bank datasets [15] (15,789 images). The expansive PMC-OA dataset encapsulates over 1.6 million image-text pairs drawn from 25 million scientific literature images and titles. Rigorous screening and preprocessing efforts are applied to distill this dataset down to 623,816 image-text pairs, with detailed steps outlined in the Data Preprocessing section. Throughout the pre-training process, we just employ text-generation supervised learning as the pre-training objective. This kind of strategy aims to establish robust representation connections between images and their corresponding texts. In the context of dialogue, the answer section is represented by captions from the text-image pairs. We have constructed 20 questions for the question section, which are categorized into two types: indicative long answers (10 sentences) and short answers (10 sentences). Depending on the length of the dialogue answer, a sentence from the relevant question type is selected as the interrogative component of the dialogue. The comprehensive list of the 20 constructed questions is provided in the appendix for reference.
MMFundus Dataset for fine-tuning VisionUnite
To realize open-ended multi-disease diagnosis of fundus images using VisionUnite and deliver comprehensive clinical explanations grounded in multiple rounds of user feedback, we curate the pioneering multimodal fundus dataset, MMFundus. MMFundus dataset comprises 296,379 pairs of fundus images and corresponding text, accompanied by 889,137 rounds of dialogue, strategically designed to enhance the instruction-following proficiency of large language models. Drawing from 52 public datasets and one private dataset, the image data in MMFundus embodies a diverse spectrum of ophthalmic conditions. The textual and dialogue components are sourced through doctor annotations, inherent category labels, and the automated generation capabilities of large language models such as InternLM [31] and GPT-4 [32]. For the testset, all image descriptions and dialogue data have undergone meticulous annotation or confirmation by medical professionals. In constructing the training set, we initially utilized doctor annotations and category labels to establish a prototype dataset. Leveraging the prototype dataset alongside a large language model, we automatically generate the image-related descriptions and dialogues batch by batch. A rigorous quality control process ensues, involving the random selection of 5% of image-text pairs and corresponding dialogues for doctor confirmation and review in the current batch. The iterative process corrects any inaccuracies in image descriptions and dialogues, ensuring that regenerated content reaches a remarkable accuracy rate of 99% in the extracted data samples. Furthermore, we design a hierarchical approach to generate descriptions at three levels—Normal/Abnormal, disease name, and clinical explanation based on the disease within the dataset. These descriptions are concatenated to form comprehensive image explanations, enriching the dataset with nuanced and clinically relevant insights. As illustrated in Figure 2, the annotation "Abnormal" signifies that the image deviates from normalcy, while the descriptor "Severe Diabetic Macular Edema" denotes the specific disease depicted in the image. Further elaborating on the fundus sign, the statement "lots of hard exudates near the macula center observed, Moderate Non-Proliferative Diabetic Retinopathy, Retinal hemorrhages, or hard exudates observed" offers a detailed explanation of this disease. The inclusion of sign-level explanations is pivotal for accurate disease diagnosis. To innovate within this realm, we introduce corresponding sign-level labels to each image, namely Vascular, Macular, FBC (Fundus Boundary Color), OCD (Optic Cup Disc), FHE (Fundus Hemorrhages Exudation), and Other. Specifically, vascular pertains to ocular artery and vein-related conditions, including retinal artery and venous occlusions, often indicated by abnormalities in the retinal arterial-venous ratio. Macular addresses diseases within or associated with the macular area, spanning macular edema and age-related macular degeneration. FBC involves signs affecting the boundary areas of diverse fundus image components and the overall fundus background, such as the Leopard Fundus generated by Pathology Myopia and extracellular deposits between the Retinal Pigment Epithelium (RPE) and Bruch’s membrane. OCD encompasses diseases related to the cup and disc, exhibiting color changes in the cup and disc area and abnormal cup-to-disc ratios, including glaucoma and associated conditions like Optical Disc Pit Maculopathy. FHE focuses on diseases related to retinal hemorrhage and exudation, encompassing conditions like Diabetic Retinopathy marked by signs like retinal hemorrhage or exudation, and Cotton wool spots, which manifest as spotted lesions on the retina. The distinction sets the MMFundus dataset apart from other multimodal biomedical datasets. By incorporating sign-level labels, we establish a three-level label system for the lower image, encompassing normal/abnormal, signs, and diseases. This nuanced system enables VisionUnite to address patient queries with granular precision, providing more comprehensive and reliable clinical explanations. Finally, details regarding the automatic generation of dialogue data with InternLM [31] and GPT-4 [32] are provided in the appendix.
Data preprocessing
In the curation and refinement of pre-training datasets, we initially gather over 1.6 million image-text pairs from the PMC-OA [13, 14] and Retina Image Bank datasets [15]. However, the PMC-OA dataset contained numerous low-quality non-biomedical images, including instances such as table images. Therefore, a meticulous screening and processing approach is undertaken. A ResNet18-based modality classifier is trained to discern non-biomedical images, with modalities defined as CT (Computed Tomography), FA (Fluorescein Angiography), Fundus, MRI (Magnetic Resonance Imaging), OCT (Optical Coherence Tomography), Pathology, PET (Positron Emission Tomography), X-ray, and Table Chart. Leveraging the predicted classification results, we retain images identified as CT, FA, Fundus, MRI, OCT, Pathology, PET, and X-ray, while discarding those with lower classification confidence to mitigate the impact of low-quality content. Concurrently, we refine the text component by removing words such as arrow, line, star, and various colors (red, yellow, blue, orange, green, purple, violet, black, white, and gray) to minimize interference from irrelevant information. Furthermore, we introduce explicit modality indicators in the image captions in response to the predicted classification outcomes, such as "This is a Fundus image." This kind of deliberate inclusion of clear modality indicators aims to equip VisionUnite with the capability to analyze modalities during the pre-training stage.
To augment the fine-tuning dataset (MMFundus), we employ the ImageEnhance Contrast method to adjust the contrast of images. It involves creating an enhancement object and utilizing the enhancement method to yield an enhanced image. The enhancement factor (a floating-point parameter) governs the degree of enhancement: an enhancement factor of 0.0 generates a pure gray image, 1.0 preserves the original image, and higher values intensify image contrast. In our model, we set the enhancement factor to 1.3. Additionally, we perform a color space adjustment, transitioning from RGB to HSV. The HSV space is better suited for image processing as it segregates color and brightness information, thereby mitigating the impact of shadows and other extraneous factors.
Baseline methods

### v2

@@ III-C The architecture of VisionUnite
We propose a large vision-language foundation model with clinical knowledge to enable the model to achieve open-ended prediction and have user interaction and clinical interpretation capabilities. Illustrated in Figure 2 (a), our proposed VisionUnite comprises a Transformer-based vision encoder, a vision adapter tailored for signs-level classification, a vision projector for aligning visual embeddings and text tokens, and a substantial language model fine-tuned on llama-7B. The large language model differs from the vision model in that it can achieve open-ended prediction. By leveraging the combination of image-text pairs and extensive dialogue information, VisionUnite excels in responding to user questions based on visual data. We introduce visual features with sign-level features into the model, which improve its ability to understand images and further enhance its clinical interpretation ability. This design can extend its ability to analyze fundus images for medical diagnosis. Our model is also designed for multi-round dialogues, which helps the model follow user instructions and achieve better problem-understanding ability.
III-C1 Transformer-Based Vision Encoder
We leverage the EVA02 model [24] with the CLIP Vision Encoder to collectively serve as the Transformer-Based Vision Encoder. Our configuration incorporates 12 layers of EvaBlock within the original vision Transformer Block [25]. Notably, in our design, the GELU activation [26] in the original vision Transformer Block is replaced with the more advanced SwiGLU [27]. It is used to enhance the performance of the FFN (position feedforward network) layer in the Transformer architecture [10]. Specifically, images, text descriptions, and questions are fed into corresponding encoders to form visual embeddings, text embeddings, and dialogue tokens during the training phase. In this phase, we use contrastive learning to supervise the encoding formation of visual and text embeddings. The VisionUnite model uses a vision encoder to extract visual embeddings which are then used by a vision adapter to make classification predictions for specific signs. This process enhances diagnostic accuracy and efficiency by identifying and localizing signs like lesions, which helps narrow down diagnostic options and leads to more focused investigations. The model reduces the need for broad differential diagnoses and increases accuracy by concentrating on clinical signs.
III-C2 Vision Adapter
The vision adapter in VisionUnite acts as a secondary encoder that takes the visual embeddings from the vision encoder and performs detailed classification into six sign category embeddings. These categories—Vascular, Macular, FBC, OCD, FHE, and Other—cover essential aspects of eye anatomy and pathology, directly correlating with common diagnostic criteria used in ophthalmology. Specifically, Vascular for blood flow, Macular for central vision, FBC for boundary conditions, OCD for optic nerve health (glaucoma), FHE for retinal bleeding and exudates, and Other for miscellaneous conditions. Then we concatenate the visual embeddings and sign category embeddings as inputs to the vision projector, which is shown in the following formulas.
$V_{embed}^{{}^{\prime}}=[V_{embed},CLS_{1}(V_{embed}),...,CLS_{6}(V_{embed})]$ (1)
$Token_{CLS}=\sum_{i=1}^{6}Token(\mathbb{I}(CLS_{i}(V_{embed})>0)K_{i})$ (2)
where $V_{embed}^{{}^{\prime}}$ represents the combination of visual embedding $V_{embed}$ and each sign category embedding $CLS_{i}(V_{embed})$. $\mathbb{I}$ is the indicator function, which can help embed the corresponding keyword embedding ($K_{i}$) and concatenate them to form $Token_{CLS}$ based on the prediction of the signs $(CLS_{i}(V_{embed})$. It demonstrates that VisionUnite is designed to mimic the diagnostic approach of medical professionals by first identifying specific signs and then narrowing down to precise disease identification. It allows VisionUnite to align with human medical expert processes, enhancing diagnostics. The sign-level classification provides an interpretable intermediate layer between raw visual features and final diagnoses too.
III-C3 Vision Projector
The vision projector is designed to integrate and synchronize the visual embeddings with dialogue tokens, facilitating a seamless multimodal interaction within the model. This component takes the concatenated visual and sign category embeddings and aligns them with the dialogue tokens by matching their feature dimensions. The inclusion of the $CLS$ token encapsulates the predicted sign-level features, which are crucial for generating relevant and context-aware text responses in the model, as shown in the formulas.
$V_{proj}=Prefix_{q}+Unsqueeze(V_{embed}^{{}^{\prime}})$ (3)
$F_{mm}=Attn(V_{proj})+[Token_{CLS},Token_{D}]$ (4)
where $V_{proj}$ represents the projected feature of enhanced visual embedding $V_{embed}^{{}^{\prime}}$. $Prefix_{q}$ represents the query prefix. The multimodal features $F_{mm}$ combine attention visual embeddings, $CLS$ tokens ($Token_{CLS}$), and dialogue tokens ($Token_{D}$). $F_{mm}$ are the inputs for training the final text generation response within the fine-tuned LLaMA model. VisionUnite integrates multiple modalities to improve interaction and understanding in multi-round dialogues, enhancing its accuracy and relevance in responding to user queries. This integration boosts its diagnostic performance and utility in clinical settings, making it a valuable tool for medical image analysis crucial for accurate diagnosis and treatment planning.
III-C4 Fine-tuned LLaMA Model
The Fine-tuned LLaMA Model is an enhanced language model based on the LLaMA-7B framework. The model has been specifically fine-tuned to integrate seamlessly with vision components, enabling advanced text generation, classification, and open-ended predictive capabilities. The aligned visual-text data is fed into the fine-tuned LLaMA model, which uses the input to generate textual output that is contextually relevant to the visual input. Unlike typical vision models, the fine-tuned LLaMA model supports interactive dialogues and dynamic user interactions. The multi-round dialogue capability allows clinicians to interrogate the reasoning process.
III-D The training objectives of VisionUnite
@@
In the architecture of VisionUnite, we have crafted three distinct training objectives to enhance convergence: image-text contrastive learning, classification supervised learning, and text-generation supervised learning. The utilization of image-text contrastive learning facilitates the refinement of visual encoders, aiding them in more effectively aligning fundus image features. Meanwhile, the application of classification-supervised learning contributes by furnishing accurate feature categories for both visual embeddings and dialogue tokens. We utilize the accuracy to guide the model training process, enhancing its overall performance. Finally, text-generation supervised learning plays a role in guiding the output of the language model, which is pivotal for achieving accurate and open-ended disease diagnoses. Unlike previous vision models, large language models are trained on vast amounts of diverse textual data. The extensive textual data enables them to generate context-aware responses and detailed explanations that are coherent and clinically relevant. We can achieve open-ended disease diagnosis using the generated text from language models. In contrast, the prediction categories of vision models are limited and concentrated on specific diseases.
III-D1 Image-text contrastive learning
@@
To attain seamless alignment between image and text features, we employ image-text contrastive learning. We utilize LLaMA [11] SentencePiece tokenizer to get the text embedding, which ensures consistency with other training objectives. We leverage the CLIP loss to quantify the similarity between image embedding and text embedding, as illustrated in Figure 2 (b) and below:
@@
III-D2 Classification supervised learning
In facilitating the acquisition of sign-level features, we employ conventional classification learning to guide the training of VisionUnite. Notably, we embrace multi-label classification while acknowledging that each sample may encompass more than one predicted category. This approach aligns more closely with the intricacies of real-world clinical scenarios, where patients may concurrently experience multiple types of eye diseases. For each category under supervision, we apply cross-entropy loss, aggregating these individual losses to derive the ultimate classification loss, as depicted in Figure 2 (c) and below:
$L_{CLS}$ $=\sum_{k=1}^{M}{L_{CLS,k}}$ (8)
$\begin{aligned} {L_{CLS,k}}&=-\frac{1}{N}{\sum_{i=1}^{N}y_{k,i}\cdot log(p_{k,i})}\\ &-\frac{1}{N}{\sum_{i=1}^{N}}{(1-y_{k,i})\cdot log(1-p_{k,i})}\end{aligned}$ (9)
where $M$ and $N$ denote the number of categories and samples respectively. $L_{CLS,k}$ signifies the cross-entropy loss for category $K$. And $y_{k,i}$ denotes whether sample $i$ belongs to class $K$, and $p_{k,i}$ represents the probability of sample $i$ in class $K$.
III-D3 Text-generation supervised learning
Within the framework of VisionUnite, we employ text-generation supervised learning to guide the text output of LLM, which is an essential aspect given that the generated text must intuitively articulate the diagnostic results and their underlying rationale. The LLM loss aims to train the model to generate text closely resembling patterns in its training data as shown in Figure 2 (d), and its formulation is articulated as follows:
@@ III-D3 Text-generation supervised learning
IV Experiments
IV-A Experimental Settings
IV-A1 The image source of MMFundus dataset
The large-scale multimodal fundus dataset (MMFundus) is constructed with images from 53 datasets. The private dataset is from a hospital in China mainland, which includes two major diseases “myopia” and “glaucoma”. The distribution is in Table. I.
TABLE I: The data distribution of MMFundus dataset
Dataset Num Dataset Num Dataset Num
HRF [28] 45 INSPIRE-AVR [29] 40 IOSTAR [30] 30
RITE [31] 40 G1020 [32] 1020 GAMMA [33] 100
ORIGA [34] 650 REFUGE [21] 1200 ODIR [18] 7000
PALM [35] 1200 RFMiD [36] 3200 RFMiDv2 [37] 860
APTOS [19] 3662 DeepDRiD [20] 1600 EyePACS [38] 35126
IDRID [39] 516 ADAM [40] 1200 ACRIMA [41] 705
MESSIDOR-2 [42] 1748 JSIEC [43] 1000 DeepEyeNet [44] 15708
LAG [45] 4854 PARAGUAY [46] 757 PAPILA [47] 488
STARE [48] 397 FIVES [49] 800 FUND [50] 179
E-ophta [51] 463 BRSET [52] 16266 MuReD [53] 2208
OIA-DDR [54] 12522 SUSTech-SYSU [55] 1219 Cataract [56] 601
DGOCF [57] 9939 BoVW [58] 2013 HarvardGlaucoma [59] 1544
RIM-ONE [60] 485 CHAKSU [61] 1345 DiaRetDB [62] 89
LSD [63] 175 GNG [64] 400 AOD [65] 14813
DHRF [66] 2757 VietAI [67] 3435 ToxoFundus [68] 411
Papilledema [69] 1369 BEH [70] 634 WilliamHoyt [71] 856
ROI [72] 1120 ROD [73] 281 BiDR [74] 2838
AIROGS [75] 101442 Private 33029
Summary 296,379 images
IV-A2 Baseline methods